# Deep Learning of Multivariate Extremes via a Geometric Representation

### v1

@@ 3.5 Assessing goodness-of-fit for DeepGauge model fits
We consider the diagnostic proposed by Wadsworth and Campbell, (2024) to assess the validity of the modelling assumption described in equation (7). Here, the fitted truncated gamma model is used to transform all of the threshold-exceeding observations of $R\mid R>r_{\tau}(\bm{w})$ to a unit exponential scale, and we compare the observed and theoretical quantiles via a QQ plot. This provides a global diagnostic for the model fit over the entire angular domain.
@@
We adapt the diagnostic proposed by Murphy-Barltrop et al., 2024b to assess goodness-of-fit with respect to the estimated extended ADF. Note that equation (6) can be reformulated as follows: given any angle $\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$, we have that ${T^{u}_{\bm{w}}\sim\text{Exp}(\Lambda(\bm{w}))}$ as $u\to\infty$, for exceedances $T^{u}_{\bm{w}}:=(T_{\bm{w}}-u)\mid(T_{\bm{w}}>u)$ with $T_{\bm{w}}$ defined in Section 3.2. Therefore, for sufficiently large $u$, the random variable $E:=-\log[1-(1-\exp\{-\Lambda(\bm{w})T^{u}_{\bm{w}}\})]=\Lambda(\bm{w})T^{u}_{\bm{w}}$ follows a unit exponential distribution, independent of the choice of $\bm{w}$. Murphy-Barltrop et al., 2024b exploit this fact and compute min-projection exceedances over the angular observations, before transforming these exceedances to the standard exponential scale. One can then evaluate the alignment of the exponential quantiles on a QQ plot. This results in a global diagnostic and allows us to test the validity of the modelling assumption described in equation (6). The corresponding algorithm for computing the diagnostic is given in Appendix B.9.
@@
Assuming unbiased estimation, the total law of probability implies that $\Pr(\bm{X}\in\mathcal{B}_{p})=p$ (Papastathopoulos et al.,, 2024). To assess the accuracy of return level set estimates, we propose the following diagnostic. First, observe that for any observation $\bm{x}_{j}\neq\bm{0}_{d}$, we have $\bm{x}_{j}\in\mathcal{B}_{p}\iff||\bm{x}_{j}||\leq r_{p}(\bm{x}_{j}/||\bm{x}_{j}||)$. Consequently, for a sample of non-zero observations $\{\bm{x}_{j}\}^{n}_{j=1}$, an empirical estimate $\hat{p}$ of $\Pr(\bm{X}\in\mathcal{B}_{p})$ is $\hat{p}:=(1/n)\sum_{j=1}^{n}\mathbbm{1}(||\bm{x}_{j}||\leq r_{p}(\bm{x}_{j}/||\bm{x}_{j}||))$. Uncertainty intervals for $\hat{p}$ can be obtained through bootstrapping the data sample. Our diagnostic is then to plot the pairs $(-\log(1-p),-\log(1-\hat{p}))$, alongside tolerance bounds, for a subset of increasing probabilities close to 1. This approach for assessing goodness-of-fit provides a multivariate extension of ‘return level plots’; see, e.g., Coles, (2001).
@@
For $d\leq 3$, we can plot the scaled sample clouds $\{\bm{x}_{j}/\log(n/2):j=1,\ldots,n\}$ against the estimated unit-level and extended ADF sets; see, e.g., Figure 2. For large enough $n$, we would expect the scaled sample clouds to lie approximately within the interiors of the estimated unit-level and extended ADF sets. Plotting scaled observations against the estimated unit-level sets has also been used for validation by, e.g., Majumder et al., 2024b ; Papastathopoulos et al., (2024). In addition to being a visual indicator of how well the estimated shapes capture the complex extremal dependence features of data, one can also use these plots to verify that the estimated unit-level sets satisfy all of the theoretical properties discussed in Section 1.
@@ 4.1 Overview
We conduct two simulation studies to investigate the efficacy of the DeepGauge framework for estimating the gauge function of three known copulas: Gaussian, Student-t, and logistic. Details of the copulas and their theoretical gauge functions are provided in Section 4.2.
Two studies are performed and their results presented in Section 4.3. Efficacy in both studies is quantified using the validation diagnostics described in Section 4.2. The first study considers, for a fixed quantile level $\tau$ and architecture, the effect of varying $d$ and $n$ on the accuracy of DeepGauge estimates of the gauge functions and exceedance probabilities; the second study considers, for a fixed dimension $d$ and sample size $n$, the effect of hyper-parameter choice. In both studies, we perform $100$ experiments for every model specification. That is, for a single choice of $n$, $d$, copula, and hyper-parameter configuration, we simulate 100 data sets and apply the methodology, separately, to each. Performance metrics are then presented as the median and $95\%$ confidence intervals over all experiments. All models for both $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ are trained over 500 iterations with a mini-batch size of 1024, and the ADF function is estimated with $u$ in (11) taken to be the $q=0.9995$ empirical quantile of the structure variable $T_{\bm{w}}$. Early-stopping is used with a patience of $\Delta=5$; see Appendix A.2 for details.
@@ 4.2 Models and performance metrics
We consider three copulas with known gauge functions and limit sets: Gaussian, Student-t, and logistic. Their theoretical unit-level sets and gauge functions are provided by Papastathopoulos et al., (2024). Recall that we consider a random vector $\bm{X}$ with standard Laplace margins. If $\bm{X}$ also has a $d$-variate Gaussian copula with positive-definite precision matrix $Q\in\mathbb{R}^{d\times d}$, its gauge function is $g(\bm{x})=\left\{\text{sgn}(\bm{x})|\bm{x}|^{1/2}\right\}^{T}Q\left\{\text{sgn}(\bm{x})|\bm{x}|^{-1/2}\right\}$ where sgn is applied component-wise. If $\bm{X}$ follows a $d$-variate Student-t copula with $Q$ as before and degrees of freedom $\nu>0$, its gauge function is $g(\bm{x})=-\nu^{-1}\sum^{d}_{i=1}|x_{i}|+(1+d\nu^{-1})\max_{i=1,\dots,d}|x_{i}|$. Finally, let $\bm{X}$ follow a logistic copula, i.e., a Gumbel copula with dependence parameter $1/\theta$ for $\theta\in(0,1]$. In this case, the form of the theoretical gauge function is cumbersome; see Appendix A.4 of Papastathopoulos et al., (2024). For the Gaussian and Student-t copulas, we randomly generate, for each $d$, a valid precision matrix. However, we impose that the corresponding correlation matrices are ordered with respect to $d$. That is, if we consider two copulas with dimension $d_{1}$ and $d_{2}$, satisfying $d_{1}<d_{2}$, then the $d_{1}\times d_{1}$ correlation matrix is a submatrix of the $d_{2}\times d_{2}$ correlation matrix. Note that the generated correlation matrices are kept fixed throughout all studies, and we further set $\nu=1$ and $\theta=0.3$
We simulate from these copulas and apply the methodology described in Section 3 to estimate the gauge function on $\mathcal{S}^{d-1}$. To quantify the accuracy of the estimated unit-level set, we measure the integrated squared error (ISE) of the form
$\mbox{ISE}=\int_{\mathcal{S}^{d-1}}\{1/g_{0}(\bm{w})-1/\tilde{g}(\bm{w})\}^{2}\mathrm{d}\bm{w},$ (12)
where $g_{0}$ and $\tilde{g}$ are, respectively, the theoretical and estimated (rescaled) gauge functions. We evaluate the integral in (12) using Monte-Carlo methods; specifically, we use the estimator $\widehat{\mbox{ISE}}=(A_{d}/|\mathcal{W}|)\sum_{\bm{w}\in\mathcal{W}}\{1/g_{0}(\bm{w})-1/\tilde{g}(\bm{w})\}^{2}$, where $A_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of $\mathcal{S}^{d-1}$, $\mathcal{W}$ denotes the set of points described in Section 3.3, and $\Gamma(\cdot)$ is the gamma function.
@@ 4.3 Results
We first consider a setting with the quantile level fixed to $\tau=0.75$ and the architecture for $g(\cdot)$ (equivalently, $\tilde{g}$) taken to be to a neural network with $N=3$ hidden layers of width 64 (see (A.1) for details), but allow the sample size $n$ and dimension $d$ to vary: we consider $d=3$, $d=5$, and $d=8$, as well as $n=10,000$, $n=50,000$, $n=100,000$, and $n=250,000$. For $r_{\tau}(\cdot)$, we use a neural network with $N=3$ hidden layers of width $32$ and keep this fixed throughout. Results for this setting are provided in Figure A2 of Appendix C. As expected, we generally observe decreasing ISE and MALE with increasing sample size $n$ across all dimensions and copulas.
@@
Figure 6 illustrates the simulation results for varying quantile level, $\tau$. We find that the optimal value of $\tau$ (in terms of minimising the performance metrics) differs across the copulas and dimension. For the Gaussian and Student-t copulas, we find that higher values of $\tau$ are preferable when estimating the unit-level set, as the ISE is reduced; the converse holds for the logistic copula. When considering the MALE, we find that the Gaussian and Student-t copulas tend to favour a lower value of $\tau$, particularly for the high dimensional case where $d=8$. Conversely, for the logistic copula, the MALE is minimised with the largest value of $\tau$. We note that the logistic copula has the quickest rate of convergence to the truncated gamma model described in (7), which may explain why this copula benefits from a smaller $\tau$ (i.e., more data is available for inference).
Table A1 in Appendix C presents the simulation results for a fixed quantile level ${\tau=0.75}$, but with varying architecture for $g(\cdot)$. We find that, for the Gaussian and Student-t copulas, the ISE is minimised when using the more complicated architecture, e.g., 3–4 layers of width 64; the converse holds for the logistic copula. Note that the logistic copula is determined by a single parameter, and so its corresponding gauge function has a much simpler shape that the other copulas (see Figure 2 for the case when $d=2$). Hence, it is unsurprising that the logistic copulas favours a simpler neural network when considering the ISE. For the MALE, the results are less clear: all three copulas generally favour an architecture with fewer parameters (and, hence, a larger effective sample size), except the Gaussian copula with $d=5$. In applications, the shape of the gauge function that arises from the data generating process is likely to be quite complex. Figure 6 and Table A1 suggest that, if the goal is to accurately estimate the unit-level set $\partial\mathcal{G}$, then one should consider using a more complicated neural network architecture for $g(\cdot)$. However, a more conservative approach should be taken if we wish to accurately estimate tail probabilities; reducing the quantile level $\tau$ and using a simpler neural network architecture will increase the effective sample size used in inference for model (7), and this improve estimation of tail probabilities.
@@
Generally, we find some variation in model fits across different choices of architecture and quantile level $\tau$, but the results are broadly similar. From this study, we can conclude that there is no one-size-fits-all approach to fully optimising hyper-parameters of the the DeepGauge framework. In practice, where the true gauge function and underlying data generating process are unknown, we advocate that practitioners fit a collection of DeepGauge models with various quantile levels and architecture, and choose the best model fit using the goodness-of-fit metrics detailed in Section 3.5.
@@ 5.1 Overview
We demonstrate the efficacy of the DeepGauge framework by modelling extremal dependence of metocean variables associated with severe ocean events. Such events typically occur when multiple metocean variables are simultaneously extreme, and pose a risk to offshore and coastal structures, such as wind farms and oil platforms (Shooter et al.,, 2021, 2022). Robust analyses of joint extremes is therefore crucial in this setting for informed decision making.
@@
Our study uses the NORA10 hindcast gridded data set (NOrwegian ReAnalysis 10km, Reistad et al.,, 2011), which provides three-hourly wave fields over the Norwegian Sea, the North Sea, and the Barents Sea, at a spatial resolution of 10km. We focus on an area between the British Isles and Iceland for the period between September 1957 and December 2009 inclusive, which amounts to $n=152,917$ observations. These data have also previously been analysed by Shooter et al., (2022). We run two analyses to illustrate the efficacy of the DeepGauge framework for dimensions $d=3$ and $d=5$. For both analyses, hyper-parameters (e.g., model architecture and quantile level $\tau$) were optimised with respect to the goodness-of-fit diagnostics described in Section 3.5. As our interest is in modelling extremal dependence and not marginal extremes, the data are apriori transformed to standard Laplace margins using a standard rank transform, applied independently at each spatial location and for each variable. All study locations are plotted in Figure 7.
@@
First, we consider the joint behaviour of wind speed (ws; m/s), significant wave height (hs; m), and mean sea level pressure (mslp; hPa) separately at three locations (01, 46, and 85, from south to north), outlined in blue in Figure 7. These three variables are associated with severe ocean events (Ewans and Jonathan,, 2014), and we aim to model their joint tail behaviour separately at each of the three locations. Of the $d=3$ variables, ws and hs exhibit positive dependence, while the other two pairs of variables exhibit negative dependence; see Figure 9. Given the complex dependence structure of these data, it is likely that standard parametric extremal dependence models will perform poorly here; hence, these data provide the perfect candidate for illustrating the flexibility of our DeepGauge model.
@@
Our second analysis models the joint tail behaviour of significant wave height (hs) across five locations (16, 33, 58, 72, and 92; highlighted in black, Figure 7) along a transect that runs approximately south-west to north-east, representing storm fronts that move along that direction. Simultaneous extremes at multiple locations pose higher risks than extremes at individual locations, and so quantification of joint extreme risk across these $d=5$ locations represents an important area for investigation.
For the $d=3$ and $d=5$ cases, we use $\tau=0.9$ and $\tau=0.4$, respectively. The neural networks for $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ in the $d=3$ setting have identical architectures: three hidden layers, each with 64 neurons. For $d=5$, both neural networks have two hidden layers; the neural networks for $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ have all widths 16 and 64, respectively.

### v2

@@ 3.5 Assessing goodness-of-fit for DeepGauge model fits
We consider the diagnostic proposed by Wadsworth and Campbell, (2024) to assess the validity of the modelling assumption described in equation (7). Here, the fitted truncated gamma model is used to transform all of the threshold-exceeding observations, i.e., $R\mid(\bm{W}=\bm{w},R>r_{\tau}(\bm{w}))$, to a unit exponential scale, and we compare the observed and theoretical quantiles via a QQ plot. This provides a global diagnostic for the model fit over the entire angular domain.
@@
We adapt the diagnostic proposed by Murphy-Barltrop et al., 2024b to assess goodness-of-fit with respect to the estimated extended ADF. Note that equation (6) can be reformulated as follows: given any angle $\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}$, we have that ${T^{u}_{\bm{w}}\sim\text{Exp}(\Lambda(\bm{w}))}$ as $u\to\infty$ for exceedances $T^{u}_{\bm{w}}:=(T_{\bm{w}}-u)\mid(T_{\bm{w}}>u)$, with $T_{\bm{w}}$ defined in Section 3.2. Therefore, for sufficiently large $u$, the random variable $E:=-\log[1-(1-\exp\{-\Lambda(\bm{w})T^{u}_{\bm{w}}\})]=\Lambda(\bm{w})T^{u}_{\bm{w}}$ approximately follows a unit exponential distribution, independent of the choice of $\bm{w}$. Murphy-Barltrop et al., 2024b exploit this fact and compute min-projection exceedances for a fixed angle, before transforming these exceedances to the standard exponential scale. One can then evaluate the alignment of the exponential quantiles on a QQ plot. This results in a local diagnostic and allows us to test the validity of the modelling assumption described in equation (6). The corresponding algorithm for computing the diagnostic is given in Appendix B.9.
@@
Assuming unbiased estimation, the total law of probability implies that $\Pr(\bm{X}\in\mathcal{B}_{p})=p$ (Papastathopoulos et al.,, 2024). To assess the accuracy of return level set estimates, we propose the following diagnostic. First, observe that for any observation $\bm{x}_{j}\neq\bm{0}_{d}$, we have $\bm{x}_{j}\in\mathcal{B}_{p}\iff||\bm{x}_{j}||\leq r_{p}(\bm{x}_{j}/||\bm{x}_{j}||)$. Consequently, for a sample of non-zero observations $\{\bm{x}_{j}\}^{n}_{j=1}$, an empirical estimate $\hat{p}$ of $\Pr(\bm{X}\in\mathcal{B}_{p})$ is $\hat{p}:=(1/n)\sum_{j=1}^{n}\mathbbm{1}(||\bm{x}_{j}||\leq r_{p}(\bm{x}_{j}/||\bm{x}_{j}||))$. Our diagnostic then plots the pairs $(-\log(1-p),-\log(1-\hat{p}))$, alongside tolerance bounds, for a subset of increasing probabilities close to 1. This approach for assessing goodness-of-fit provides a multivariate extension of ‘return level plots’; see, e.g., Coles, (2001).
@@
For $d\leq 3$, we can plot the scaled sample clouds $\{\bm{x}_{j}/\log(n/2):j=1,\ldots,n\}$ against the estimated unit-level and extended ADF sets; see, e.g., Figure 2. For large enough $n$, we would expect the scaled sample clouds to lie approximately within the interiors of the estimated unit-level and extended ADF sets. Plotting scaled observations against the estimated unit-level sets has also been used for validation by, e.g., Majumder et al., 2024b ; Papastathopoulos et al., (2024). In addition to being a visual indicator of how well the estimated shapes capture the complex extremal dependence features of data, one can also use these plots to verify that the estimated unit-level and extended ADF sets satisfy all of the theoretical properties discussed in Sections 1 and 2.
@@ 4.1 Overview
We conduct two simulation studies to investigate the efficacy of the DeepGauge framework for estimating the gauge functions on $\mathcal{S}^{d-1}$ (equivalently, the unit-level sets) of three known copulas: Gaussian, Student-t, and logistic. Details of the copulas and their theoretical gauge functions are provided in Section 4.2.
Two studies are performed and their results presented in Section 4.3. Efficacy in both studies is quantified using the validation diagnostics described in Section 4.2. The first study considers, for a fixed quantile level $\tau$ and architecture, the effect of varying $d$ and $n$ on the accuracy of DeepGauge estimates of the gauge functions on $\mathcal{S}^{d-1}$ and exceedance probabilities. The second study considers, for a fixed dimension $d$ and sample size $n$, the effect of hyper-parameter choice. In both studies, we perform $100$ experiments for every model specification. That is, for a single choice of $n$, $d$, copula, and hyper-parameter configuration, we simulate 100 data sets and apply the methodology, separately, to each. Performance metrics are then presented as the median and $95\%$ confidence intervals over all experiments. All models for both $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ are trained over 500 iterations with a mini-batch size of 1024, and the ADF function is estimated with $u$ in (11) taken to be the $q=0.9995$ empirical quantile of the structure variable $T_{\bm{w}}$. Early-stopping is used with a patience of $\Delta=5$; see Appendix A.2 for details.
@@ 4.2 Models and performance metrics
We consider three copulas with known gauge functions and limit sets: Gaussian, Student-t, and logistic; the theoretical unit-level sets and gauge functions associated with them are detailed in Papastathopoulos et al., (2024). If the random vector $\bm{X}$ also has a $d$-variate Gaussian copula with positive-definite precision matrix $Q\in\mathbb{R}^{d\times d}$, its gauge function is $g(\bm{x})=\left\{\text{sgn}(\bm{x})|\bm{x}|^{1/2}\right\}^{T}Q\left\{\text{sgn}(\bm{x})|\bm{x}|^{-1/2}\right\}$, where $\text{sgn}(\cdot)$ is applied component-wise. If $\bm{X}$ follows a $d$-variate Student-t copula with $Q$ as before and degrees of freedom $\nu>0$, its gauge function is $g(\bm{x})=-\nu^{-1}\sum^{d}_{i=1}|x_{i}|+(1+d\nu^{-1})\max_{i=1,\dots,d}|x_{i}|$. Finally, let $\bm{X}$ follow a logistic copula, i.e., a Gumbel copula with dependence parameter $1/\theta$ for $\theta\in(0,1]$. In this case, the form of the theoretical gauge function is cumbersome; see Appendix A.4 of Papastathopoulos et al., (2024). For the Gaussian and Student-t copulas, we randomly generate, for each $d$, a valid precision matrix. However, we impose that the corresponding correlation matrices are ordered with respect to $d$. That is, if we consider two copulas with dimension $d_{1}$ and $d_{2}$, satisfying $d_{1}<d_{2}$, then the $d_{1}\times d_{1}$ correlation matrix is a submatrix of the $d_{2}\times d_{2}$ correlation matrix. Note that the generated correlation matrices are kept fixed throughout all studies, and we further set $\nu=1$ and $\theta=0.3$
We simulate data from these copulas and apply the methodology described in Section 3 to estimate the gauge function on $\mathcal{S}^{d-1}$. To quantify the accuracy of the estimated unit-level set, we measure the integrated squared error (ISE) of the form
$\mbox{ISE}:=\int_{\mathcal{S}^{d-1}}\{1/g_{0}(\bm{w})-1/\tilde{g}(\bm{w})\}^{2}\mathrm{d}\bm{w},$ (12)
where $g_{0}$ and $\tilde{g}$ are, respectively, the theoretical and estimated gauge functions. We evaluate the integral in (12) using Monte-Carlo methods; specifically, we use the estimator $\widehat{\mbox{ISE}}:=(A_{d}/|\mathcal{W}|)\sum_{\bm{w}\in\mathcal{W}}\{1/g_{0}(\bm{w})-1/\tilde{g}(\bm{w})\}^{2}$, where $A_{d}:=2\pi^{d/2}/\Gamma(d/2)$ is the surface area of $\mathcal{S}^{d-1}$, $\mathcal{W}$ is the set of points described in Section 3.3, and $\Gamma(\cdot)$ is the gamma function.
@@ 4.3 Results
We first consider a setting with the quantile level fixed to $\tau=0.75$ and the architecture for $g(\cdot)$ (equivalently, $\tilde{g}$) taken to be to a neural network with $N=3$ hidden layers of width 64 (see Appendix A.1 for details), but allow the sample size $n$ and dimension $d$ to vary: we consider $d=3$, $d=5$, and $d=8$, as well as $n=10,000$, $n=50,000$, $n=100,000$, and $n=250,000$. For $r_{\tau}(\cdot)$, we use a neural network with $N=3$ hidden layers of width $32$ and keep this fixed throughout. Results for this setting are provided in Figure A2 of Appendix C. As expected, we generally observe decreasing ISE and MALE with increasing sample size $n$ across all dimensions and copulas.
@@
Figure 6 illustrates the simulation results for varying quantile level, $\tau$. We find that the optimal value of $\tau$ (in terms of minimising the performance metrics) differs across the copulas and dimension. For the Gaussian and Student-t copulas, we find that higher values of $\tau$ are preferable when estimating the unit-level set, as the ISE is reduced; the converse holds for the logistic copula. When considering the MALE, we find that the Gaussian and Student-t copulas tend to favour a lower value of $\tau$, particularly for the high dimensional case where $d=8$. Conversely, for the logistic copula, the MALE is minimised with the largest value of $\tau$. We note that the logistic copula has the quickest rate of convergence to the truncated gamma model described in (7) (Wadsworth and Campbell,, 2024), which may explain why this copula benefits from a smaller $\tau$.
Table A1 in Appendix C presents the results for a fixed quantile level ${\tau=0.75}$, but with varying architecture for $g(\cdot)$. We find that, for the Gaussian and Student-t copulas, the ISE is minimised when using the more complicated architecture, e.g., 3–4 layers of width 64; the converse holds for the logistic copula. Note that the logistic copula is determined by a single parameter, and so its corresponding gauge function has a much simpler shape that the other copulas (see Figure 2 for the case when $d=2$). Hence, it is unsurprising that the logistic copulas favours a simpler neural network when considering the ISE. For the MALE, the results are less clear: all three copulas generally favour an architecture with fewer parameters (and, hence, a larger effective sample size), except the Gaussian copula with $d=5$. In applications, the shape of the gauge function that arises from the data generating process is likely to be quite complex. Figure 6 and Table A1 suggest that, if the goal is to accurately estimate the unit-level set $\partial\mathcal{G}$, then one should consider using a more complicated neural network architecture for $g(\cdot)$. However, a more conservative approach should be taken if we wish to accurately estimate tail probabilities; reducing the quantile level $\tau$ and using a simpler neural network architecture will increase the effective sample size used in inference for model (7), and this improve estimation of tail probabilities.
@@
Generally, we find some variation in model fits across different choices of architecture and quantile level $\tau$, but the results are broadly similar. From this study, we can conclude that there is no one-size-fits-all approach to fully optimising hyper-parameters of the the DeepGauge framework. In practice, where the true gauge function and underlying data generating process are unknown, we advocate that practitioners fit a collection of DeepGauge models with various quantile levels and architectures, and choose the best model fit using the goodness-of-fit metrics detailed in Section 3.5.
@@ 5.1 Overview
We demonstrate the efficacy of the DeepGauge framework by modelling extremal dependence among metocean variables associated with severe ocean events. Such events typically occur when multiple metocean variables are simultaneously extreme, and pose a risk to offshore and coastal structures, such as wind farms and oil platforms (Shooter et al.,, 2021, 2022). Robust analyses of joint extremes is therefore crucial for informed decision making.
@@
Our study uses the NORA10 hindcast gridded data set (NOrwegian ReAnalysis 10km, Reistad et al.,, 2011), which provides three-hourly wave fields over the Norwegian Sea, the North Sea, and the Barents Sea, at a spatial resolution of 10km. We focus on an area between the British Isles and Iceland for the period between September 1957 and December 2009 (inclusive), which amounts to $n=152,917$ observations. These data have also previously been analysed by Shooter et al., (2022). We run two analyses to illustrate the efficacy of the DeepGauge framework for dimensions $d=3$ and $d=5$. For both analyses, hyper-parameters (e.g., model architecture and quantile level $\tau$) were optimised with respect to the goodness-of-fit diagnostics described in Section 3.5. As our interest is in modelling extremal dependence and not marginal extremes, the data are apriori transformed to standard Laplace margins using a standard rank transform, applied independently at each spatial location and for each variable. All study locations are plotted in Figure 7.
@@
First, we consider the joint behaviour of wind speed (ws; m/s), significant wave height (hs; m), and mean sea level pressure (mslp; hPa) separately at three locations (01, 46, and 85, from south to north), outlined in blue in Figure 7. These three variables are associated with severe ocean events (Ewans and Jonathan, 2014a, ), and we aim to model their joint tail behaviour separately at each of the three locations. Of the $d=3$ variables, ws and hs exhibit positive dependence, while the other two pairs of variables exhibit negative dependence; see Figure 9. Given the complex dependence structure of these data, it is likely that standard parametric extremal dependence models will perform poorly here; hence, these data provide the perfect candidate for illustrating the flexibility of our DeepGauge model.
@@
Our second analysis models the joint tail behaviour of significant wave height (hs) across five locations (16, 33, 58, 72, and 92; highlighted in black, Figure 7) along a transect that runs approximately south-west to north-east, representing storm fronts that move along that direction. Simultaneous extremes at multiple locations pose higher risks than extremes at individual locations (Simpson and Wadsworth,, 2021), and so quantification of joint extreme risk across these $d=5$ locations represents an important area for investigation.
For the $d=3$ and $d=5$ cases, we use $\tau=0.9$ and $\tau=0.4$, respectively. The neural networks for $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ in the $d=3$ setting have identical architectures: three hidden layers, each with 64 neurons. For $d=5$, both neural networks have two hidden layers; the neural networks for $r_{\tau}(\cdot)$ and $\tilde{g}(\cdot)$ have all widths 16 and 64, respectively. Since the data exhibit strong temporal dependence, uncertainty is quantified via block bootstrapping: this sampling scheme retains temporal dependence in the resampled data set, ensuring the additional uncertainty that arises due to lower effective sample sizes is accounted for (Politis and Romano,, 1994). Moreover, the block size set to $4$ days, since this length is appropriate for capturing the temporal dependence typically observed in metocean data sets (Murphy-Barltrop et al., 2024a, ).